OpenStudy (anonymous):

Find the equation of the tangent line to the curve y=sec(x)-2cos(x) at the point (pi/3,1).

OpenStudy (gorv):

first find slope of tangent that will be = dy/dx

OpenStudy (gorv):

dy/dx=d(secx-2cosx)/dx

OpenStudy (gorv):

and after finding it put these (pi/3,1) in the result of dy/dx

OpenStudy (gorv):

calculate it n tell me the ans ..i assist you in this plsss try

OpenStudy (anonymous):

y=sec(x)tan(x)-2sin(x)

OpenStudy (anonymous):

i dont get how to put this in to y=mx+b